# Edge-Cloud Collaborative Computing on Distributed Intelligence and Model Optimization: A Survey

### v2

@@ VIII-A2 System Integration Challenges
Effectively integrating the diverse hardware and software of edge-cloud collaborative systems presents significant challenges, demanding effective heterogeneity management, robust interoperability solutions, and standardized protocols [290]. Managing heterogeneous hardware and software across resource-constrained edge nodes and powerful cloud servers requires addressing variations in processing capabilities, memory capacities, network bandwidths, and operating systems [257]. Ensuring seamless communication and data exchange between these disparate components is crucial, as differing data formats, communication protocols, and software interfaces can create interoperability issues [291]. Efficient workload orchestration across heterogeneous resources necessitates sophisticated scheduling algorithms that consider each device’s capabilities and limitations [126], while application portability demands flexible deployment mechanisms. Approaches like containerization (e.g., Docker [290]) and WebAssembly  [291] offer portability solutions. AI-driven techniques [292, 293] aim to optimize task placement and resource management. Additionally, interoperability challenges arise from the heterogeneous mix of devices with distinct operating systems, communication protocols, and data formats [44]. Variations in data representation and communication protocols can lead to compatibility issues, requiring complex data transformations [293]. Standardized communication protocols and well-defined interfaces are essential for seamless data exchange, simplified application development, and enhanced performance and scalability [261, 292]. Such standardization promotes interoperability between vendor solutions, fostering a more open ecosystem [280, 294]. Standard protocols ensure consistent and reliable communication, facilitating efficient data exchange and coordinated operation across diverse components [291, 222]. Standardized data exchange formats promote interoperability, enabling tasks like distributed model training and real-time analytics [292]. Consequently, adopting standard protocols simplifies system integration, reduces development complexity, and accelerates the deployment of edge-cloud solutions [293, 261]. This is particularly critical for applications like autonomous driving [44] and smart healthcare [133] where reliable, real-time data exchange is paramount [295].
@@ VIII-A3 Performance Bottlenecks
ECCC faces significant performance bottlenecks stemming from resource limitations. Communication overhead introduces latency and bandwidth consumption between edge devices and the cloud, impacting real-time applications [98, 253, 296]. Strategies like optimized task offloading, latency-aware strategies, data compression, and efficient communication protocols can mitigate these issues [297, 55, 282, 133]. Additionally, edge devices, often constrained by limited processing power, memory, and energy, struggle to execute complex AI models and handle large datasets, especially with the growing data volume from IoT devices [57, 298, 299, 26]. This computational bottleneck is further complicated by device heterogeneity [17, 260]. Limited storage capacity at the edge, exacerbated by data volume and device heterogeneity, necessitates efficient data management solutions like intelligent caching, data compression, distributed storage (e.g., EdgeKV [54]), and data lifecycle management [39, 300, 301, 195, 133, 302]. Furthermore, real-time applications, such as autonomous driving, industrial automation, and smart healthcare, demand ultra-low latency [240, 303, 18]. Consequently, future research should focus on minimizing latency through techniques like data compression, model partitioning, and adaptive resource allocation [195, 18].
@@ VIII-B Future Research Directions
A critical challenge lies in scaling AI model deployments across expansive edge-cloud systems while preserving performance and efficiency as device numbers and data volumes surge. Promising research avenues include investigating advanced AI techniques such as neural architecture search, automated machine learning, and large language models. Additionally, exploring AI agents, embodied AI, neuromorphic and quantum edge computing, along with 6G network integration, offers potential for revolutionizing these systems. Equally crucial is improving system-level aspects, including dynamic resource management, adaptive optimization strategies, and robust fault tolerance mechanisms to enhance reliability.
@@ VIII-B1 Advanced AI Techniques
Figure 10: Overview of advanced AI techniques, including NAS, AutoML, LLMs, AI agents, embodied AI, neuromorphic computing, quantum edge computing, and 6G integration.
As illustrated in Fig. 10 shows the summary of advanced AI technologies in future research directions, including NAS, AutoML, and LLMs, contributes to enhancing system efficiency and adaptability. All these modules operate synergistically under the orchestration of an edge-cloud control unit, which facilitates dynamic task allocation, resource optimization, and intelligent decision-making across heterogeneous infrastructures.
$\bullet$ Neural Architecture Search  NAS automates the design and optimization of neural network architectures, aiming to discover models that outperform manually designed ones in accuracy and efficiency for edge-cloud collaborative environments [304]. NAS algorithms explore a vast search space of potential architectures, evaluating their performance to identify architectures tailored to the specific constraints of edge-cloud systems, such as limited resources and bandwidth [104, 248].  Approaches like reinforcement learning, evolutionary algorithms, and gradient-based NAS show promise in finding optimal architectures for specific tasks and resource constraints [49], and combining NAS with model optimization techniques like compression and adaptation further improves performance and efficiency [305, 306].  This integrated approach addresses the challenges of deploying complex AI models in heterogeneous edge-cloud environments [307].
$\bullet$ AutoML Automated machine learning (AutoML) streamlines AI model development and deployment in edge-cloud systems by automating processes like model selection, hyperparameter tuning, and deployment [308, 309]. This automation is crucial in edge-cloud environments where diverse hardware and software coupled with resource constraints make manual optimization complex [310].  For example, techniques like FAST-DAD [311] distill complex models into smaller, deployable ones for edge devices, while frameworks like H2O AutoML simplify model selection and hyperparameter optimization [309].  Language-driven systems like AutoMMLab [312] further democratize AI development.  However, challenges remain in transparency, flexibility for complex scenarios, and consistent coverage of the entire ML workflow [313]. In large-scale environments, sampling-based AutoML and human-centered systems like OmniForce [314, 310] show promise, while automatically generating data-dependent configurations improves model performance [315].
@@
$\bullet$ Embodied AI Embodied AI, focusing on agents that interact with their environment, requires new algorithms and frameworks tailored for resource-constrained edge devices [321].  Algorithms for robot control and coordination must consider latency and bandwidth limitations [322], while efficient task offloading and resource allocation optimize performance [280].  Exploring embodied AI for multi-robot collaboration and dynamic visual servoing offers significant opportunities [323, 322].  Developing robust and adaptive algorithms that leverage distributed resources is a critical research direction [324].
@@ IX Conclusion
In this survey, we comprehensively examine AI-driven edge-cloud collaborative computing, focusing on distributed intelligence and model optimization. We first outline foundational architectures, then categorize key techniques into model optimization, resource management, and privacy-security enhancements. Additionally, we highlight trends in AI-driven resource allocation and model adaptation. Despite notable progress, challenges in system heterogeneity, security, and application diversity persist, urging further research into advanced AI techniques and emerging technologies like 6G, LLM integration, AI agents, embodied AI, edge quantum computing, etc.

### v4

@@ VIII-A2 System Integration Challenges
System integration in ECCC demands sophisticated heterogeneity management across diverse hardware capabilities, operating systems, network bandwidths, and communication protocols between resource-constrained edge nodes and powerful cloud servers [257, 225]. Interoperability challenges stem from differing data formats, communication protocols, and software interfaces, necessitating complex data transformations and sophisticated scheduling algorithms that consider device-specific capabilities [258, 259, 128]. Standardized protocols and well-defined interfaces become essential for seamless data exchange, simplified development, and enhanced scalability, with containerization (Docker [257]) and WebAssembly [258] providing portability solutions alongside AI-driven task placement optimization [260, 259]. Standardization particularly benefits critical applications like autonomous driving [53] and smart healthcare [142] requiring reliable real-time data exchange, while promoting vendor interoperability and fostering open ecosystems [247, 261, 229].
@@ VIII-A3 Performance Bottlenecks
ECCC systems encounter significant performance bottlenecks from communication overhead introducing latency and bandwidth consumption that impacts real-time applications [116, 221, 262], while edge devices struggle with limited processing power, memory, and energy constraints when executing complex AI models and handling growing IoT data volumes [66, 263, 264, 29]. Resource limitations become particularly acute during test-time adaptation scenarios where models must adjust to new data distributions while maintaining strict latency and energy budgets [115]. Storage limitations exacerbated by device heterogeneity necessitate intelligent caching, data compression, distributed storage solutions like EdgeKV [63], and comprehensive data lifecycle management [48, 265, 266, 267]. Data distribution shifts across geographically distributed edge nodes create additional performance challenges, requiring adaptive load balancing and intelligent workload redistribution to maintain system efficiency. Real-time applications including autonomous driving, industrial automation, and smart healthcare demanding ultra-low latency [208, 268, 21] require integrated architectural innovations combining hierarchical edge computing with intelligent resource management [269], utilizing frameworks like EdgeMatrix [247] for multi-agent scaling, KubeEdge [195] for automated provisioning, predictive offloading [123], intelligent caching [270], and hardware acceleration technologies for optimal performance.
VIII-A4 Practical Implementation Barriers
Real-world ECCC deployment reveals substantial barriers extending beyond theoretical optimization to encompass operational complexities, economic constraints, and integration challenges, where device heterogeneity necessitates standardized middleware layers and adaptive frameworks for automatic capability detection [20, 28, 25, 60, 49]. Network connectivity issues require robust offline operation modes and intelligent synchronization mechanisms [35], while deployment complexity often results in 40-60% timeline overruns due to integration challenges and vendor compatibility issues [226]. Economic viability concerns balance substantial upfront investments against long-term benefits, with edge computing reducing bandwidth costs by 30-50% but requiring 3-5 year payback periods [142, 45], while technical pitfalls include underestimated communication overhead [21, 253], data consistency issues requiring eventual consistency models [10], and resource contention demanding sophisticated isolation algorithms [13, 29]. Successful implementations emphasize phased rollout strategies, hybrid architectures, comprehensive security frameworks, and dedicated cross-functional teams [18, 19, 14].
@@ VIII-B Future Research Directions
Future research in ECCC must address critical scalability challenges while advancing revolutionary computing paradigms [57]. Short-term objectives include developing adaptive model partitioning algorithms [58] and standardized APIs for heterogeneous device integration [200], alongside lightweight containerization technologies optimized for resource-constrained environments. Furthermore, medium-term objectives should focus on neuromorphic computing integration [271, 272] for ultra-low-power edge inference and quantum-resistant security protocols [273, 274] for long-term protection. Additionally, long-term research directions encompass quantum edge computing [275, 276] for solving intractable optimization problems and fully autonomous edge-cloud ecosystems capable of self-optimization [259]. Consequently, implementation feasibility analysis indicates that standardization efforts represent the most immediately achievable objectives, while quantum computing integration and autonomous systems present longer-term challenges requiring significant advances in underlying technologies [17].
@@ VIII-B1 Advanced AI Techniques
Figure 9: Overview of advanced AI techniques.
Advanced AI technologies represent the next frontier in ECCC, offering transformative capabilities through intelligent automation [277] and revolutionary computational paradigms. As illustrated in Fig. 9, algorithm optimization techniques including NAS [15] and AutoML [278, 279] enable autonomous system optimization by automatically discovering optimal architectures tailored to edge-cloud constraints, thereby reducing manual configuration overhead across heterogeneous infrastructure. Furthermore, intelligent interaction capabilities encompass AI agents [260] that facilitate natural human-system interaction and autonomous decision-making for adaptive responses to dynamic conditions. Additionally, hardware & infrastructure innovations particularly neuromorphic computing [271], quantum edge computing [275], and 6G network integration [26, 280] promise revolutionary advances in computational efficiency and connectivity. Consequently, the synergistic integration of these technologies enables unprecedented system autonomy, where NAS optimizes architectures for edge constraints while AI agents coordinate workflows and quantum computing tackles intractable optimization problems, collectively transforming edge-cloud system capabilities.
$\bullet$ Neural Architecture Search  NAS automates the design and optimization of neural network architectures, aiming to discover models that outperform manually designed ones in accuracy and efficiency for edge-cloud collaborative environments [281]. NAS algorithms explore a vast search space of potential architectures, evaluating their performance to identify architectures tailored to the specific constraints of edge-cloud systems, such as limited resources and bandwidth [144, 216].  Approaches like reinforcement learning, evolutionary algorithms, and gradient-based NAS show promise in finding optimal architectures for specific tasks and resource constraints [58], and combining NAS with model optimization techniques like compression and adaptation further improves performance and efficiency [282, 283].  Ultimately, integrated optimization addresses the challenges of deploying complex AI models in heterogeneous edge-cloud environments [284].
$\bullet$ AutoML Automated machine learning (AutoML) streamlines AI model development and deployment in edge-cloud systems by automating processes like model selection, hyperparameter tuning, and deployment [285, 277]. Automated optimization becomes crucial in edge-cloud environments where diverse hardware and software coupled with resource constraints make manual optimization complex [286].  For example, techniques like FAST-DAD [287] distill complex models into smaller, deployable ones for edge devices, while frameworks like H2O AutoML simplify model selection and hyperparameter optimization [277].  Language-driven systems like AutoMMLab [278] further democratize AI development.  However, challenges remain in transparency, flexibility for complex scenarios, and consistent coverage of the entire ML workflow [279]. In large-scale environments, sampling-based AutoML and human-centered systems like OmniForce [288, 286] show promise, while automatically generating data-dependent configurations improves model performance [289].
@@
$\bullet$ Embodied AI Embodied AI, focusing on agents that interact with their environment, requires new algorithms and frameworks tailored for resource-constrained edge devices [295].  Algorithms for robot control and coordination must consider latency and bandwidth limitations [296], while efficient task offloading and resource allocation optimize performance [247]. Furthermore, exploring embodied AI applications for multi-robot collaboration and dynamic visual servoing presents significant research opportunities [297, 296]. Ultimately, developing robust and adaptive algorithms that effectively leverage distributed computational resources represents a critical research direction for advancing embodied AI in edge-cloud environments [298].
@@ IX Conclusion
In this survey, we comprehensively examine AI-driven edgecloud collaborative computing, focusing on distributed intelligence and model optimization. We first outline foundational architectures, then categorize key techniques into model optimization, resource management, and privacy-security enhancements. Additionally, we highlight trends in AI-driven resource allocation and model adaptation. Despite notable progress, challenges in system heterogeneity, security, and application diversity persist, urging further research into advanced AI techniques and emerging technologies like 6G, LLM integration, AI agents, embodied AI, edge quantum computing, etc.